# Fourier PINNs: From Strong Boundary Conditions to Adaptive Fourier Bases

### v1

@@ 4 Fourier Analysis of Strong BC PINNs
In this section, we explain why the strong BC PINNs outperform standard PINNs using Fourier analysis. Both NNs and PINNs often exhibit spectral bias, quickly capturing low-frequency information but struggling with high-frequency components (Basri et al., 2020; Rahaman et al., 2019; Xu et al., 2019; Wang et al., 2021). This bias means that while NNs can quickly grasp general patterns (low frequencies), they struggle with capturing finer details (high frequencies), leading to potential inaccuracies in the solution. The frequency spectrum of a PINN showing spectral bias would show large coefficients at low frequencies, indicating that the NN has effectively captured the low-frequency components of the target solution. In contrast, the coefficients for high-frequency components would decay slowly and exhibit heavy tails, indicating that the network is not effectively filtering out the high-frequency noise or inaccuracies. These tails suggest that the high-frequency components, even though not dominant, still have significant amplitudes. Overall, the frequency spectrum would highlight a pronounced low-frequency peak with a gradual and insufficient reduction in the magnitude of higher frequencies, reflecting the network’s difficulty in learning and accurately representing the high-frequency aspects of the solution.
@@
Figure 3: The frequency spectrum of the learned solution surrogate (a) and the convolution operation in the strong BC PINN ($\phi_{poly}$) with $\omega=15$. In both cases, the ground-truth solution is $\sin(\omega x)$. The left graph shows how the standard and strong BC PINNs handle frequency components, while the right graph illustrates the impact of the convolution operation on reducing high-frequency noise.
Figure 3 shows the frequency spectrum from the discrete Fourier transform of a standard PINN trained to solve the 1D Poisson problem defined in equation 12 with the true solution $u(x)=\sin(\omega x)$ for $x\in[0,2\pi]$, where the solution frequency is $\omega=15$. Although the standard PINN identifies the target frequency seen by the large coefficient at $\omega=15$, it exhibits large coefficients for higher frequencies, leading to heavy tails and reduced accuracy. In contrast, the frequency spectrum of the strong BC PINN (Figure 3) demonstrates that it not only captures the target frequency as accurately as the standard PINN, but its higher frequency coefficients decay much faster. This rapid decay weakens or excludes the influence of unnecessary high frequencies, resulting in significantly improved accuracy. In simpler terms, while both networks can identify the main pattern in the data (the target frequency), the strong BC PINN is better at ignoring unnecessary noise (high-frequency components) which leads to a more accurate overall solution.
To analyze why the distance function ($\phi(x)$) after multiplied to the NN in equation 8 can help obtain better coefficients for high frequencies, we first represent the general boundary function $\phi(x)$ as an infinite Fourier series,
@@
where $i$ indicates the imaginary part and $n$ are integers. We first analyze the polynomial distance function by substituting equation 9 for $\hat{\phi}$ in equation 15 giving,
@@
Similarly, to compare against our proposed adaptive distance function, we substitute in equation 10 for $\hat{\phi}$ in equation 15 and obtain,
@@
Here, we expand the boundary functions $\phi(x)$ to be periodic (i.e., duplicating its definition in $[0,2\pi]$ to other intervals) and the period is $2\pi$. We can now obtain the Fourier transforms of each $\phi_{\infty}(x)$ by evaluating,
@@
Since $|\hat{\phi}_{poly}[n]|\propto 1/n^{2}$, the amplitude of the frequencies of $\phi_{poly}(x)$ decay quadratically fast with the increase of the absolute frequency value. This rapid decay is beneficial for denoising as it effectively suppresses high-frequencies leading to smoother approximations. However, it may result in the loss of fine details, especially if important signal components are present at higher frequencies. Alternatively, $|\hat{\phi}_{exp}[n]|\propto\alpha/(\alpha^{2}+n^{2})$, indicating that the frequency amplitudes follow Lorentzian decay controlled by the decay rate $\alpha$. Lorentzian decay is less aggressive in reducing high-frequency components, which helps preserve fine details and sharp transitions in the signal. This flexibility is useful in tuning the decay based on signal characteristics, however, the slower decay rate may lead to less effective noise suppression.
@@
$\hat{u}_{\bm{\theta}}(\omega)=\hat{\phi}*\widehat{u}_{N}=\int_{-\infty}^{\infty}\hat{\phi}(\omega-\nu)\;\widehat{u}_{N}(\nu)\;{\rm d}\nu$ (19)
where * denotes convolution, and $\widehat{u}_{N}$ is the Fourier transform of the NN. Suppose $\omega>0$, we know that $\hat{\phi}(\omega-\nu)$ is obtained by first reflecting $\hat{\phi}(\nu)$ about the y-axis, and then shifting the frequency $\omega$ left along the spectrum. Then, we integrate $\widehat{u}_{N}(\nu)$ weighted by $\hat{\phi}(\omega-\nu)$. The larger the frequency $\omega$, the more $\hat{\phi}$ is moved left to obtain $\hat{\phi}(\omega-\nu)$ resulting in a stronger weighting of the high-frequency components during integration. Since the frequency coefficients of $\hat{\phi}$ decay quadratically fast ($|\hat{\phi}[n]|\propto 1/n^{2}$), the larger the portion of the tail is used, and the smaller the integration result. Specifically, with $\phi$, the corresponding amplitude of $\hat{u}_{\bm{\theta}}(\omega)$ decrease fast when $\omega$ increases [^1] The same conclusion applies when we consider $\omega<0$.. See Figure 3 for an illustration. Thereby, during the course of training, the boundary function pushes the surrogate model in equation 8 to weaken irrelevant high-frequency components to reduce large tails and to better capture the frequency spectrum.
@@ 5 Fourier PINNs
Our Fourier analysis in the previous section revealed that the standard PINNs inefficiently capture the true amplitude of frequencies in the power spectrum of their predicted solutions. Specifically, the analysis showed that although standard PINNs could identify the true solution frequency, they failed to prune large and noisy frequencies, as indicated by the large amplitudes assigned at the tail end of the power spectrum in Figure 3. While the strong BC PINNs address this issue by multiplying the NN with a boundary function—which helps decay high-frequency amplitudes faster—their overall benefit remains limited. As shown in Figure 2, performance degrades as $k$ increases, indicating the boundary function’s inability to learn ultra-high frequency solutions. Additionally, the strong BC PINNs are restricted to specific boundary conditions, such as Dirichlet conditions in equation 7, limiting their applicability.
@@ 5.1 Fourier PINNs
To overcome these challenges, we introduce Fourier PINNs, a novel PINN architecture. Fourier PINNs seamlessly handle diverse boundary conditions, like the standard PINNs, while emphasizing the true solution frequencies during training—akin to the strong BC PINNs. Specifically, in order to flexibly yet comprehensively capture the frequency spectrum, we first introduce a set of dense frequency candidates, $\{\omega_{n}\}_{n=1}^{K}$, evenly sampled from the range $[1,K]$. From this set of frequency candidates, we define a set of trainable Fourier bases $u_{B}$ as,
@@
While we present our method in the one dimensional input case, it is straightforward to extend to multiple dimensions. Consider the two dimensional input ${\bf x}=[x_{1},x_{2}]$ for an example. We use the same set of frequencies to construct Fourier bases for each input. Specifically, we construct
@@
Then we apply the cross-product of $\bm{\phi}(x_{1})$ and $\bm{\phi}(x_{2})$ to obtain the tensor-product Fourier bases for ${\bf x}$. The surrogate model is given by,
@@ 5.2 Adaptive Basis Selection Training Algorithm
To aid in identifying the significant frequencies, we add an $L_{2}$ regularization term to the basis parameters ${\bf w}$. This additional regularization penalizes large coefficients for the basis functions, thereby promoting sparsity and encouraging the model to focus on learning the most significant frequencies. The full loss function we use is similar to PINNs (see equation 4), but with the addition of the $L_{2}$ regularization term. Specifically, the loss function is,
@@
where $\|{\bf w}\|^{2}$ represents the $L_{2}$ regularization of ${\bf w}$ and $\alpha$ is the regularization strength defined by the user prior to initiating training. This regularization term helps prune or inhibit unnecessary frequencies, allowing the model to focus more on learning and improving the bases that significantly contribute to the solution.
Our adaptive basis selection algorithm optimizes the regularized loss function in equation 23 by building upon the hybrid least squares gradient descent method proposed by Cyr et al. (2020). Cyr et al. (2020) formulates the output of a neural network as a linear combination of non-linear basis functions, as shown in equation 3. Their algorithm alternates between optimizing the hidden weights (${\bm{\theta}}^{H}$) via gradient descent and the final layer coefficients ($\mathbf{c}$) through the least squares problem of the form,
$\operatorname*{arg\,min}\nolimits_{\textbf{c}}\|A\textbf{c}-\textbf{b}\|_{\ell_{2}}^{2},$ (24)
such that $\textbf{b}_{i}={\mathcal{L}}[u]({\bf x}^{i})$ for the data points $\{{\bf x}^{i}\}_{i=1}^{M}$, $A_{ij}={\mathcal{L}}[\psi_{j}({\bf x}_{i};{\bm{\theta}}^{H})]$ for $j=1,...,W$, and ${\mathcal{L}}$ is some linear operator. equation 24 extends to problems with multiple operators, such as those defined by equation 1 and equation 2 when ${\mathcal{F}}$ and ${\mathcal{B}}$ are linear operators.
Representing the output of Fourier PINNs in equation 20 as a linear combination of both adaptive bases (i.e., the hidden layers of the NN) and an augmented Fourier series ensures our model’s compatibility with the general approach of the alternating least squares and gradient descent optimization. While Cyr et al. (2020) hold the coefficients c constant while optimizing the bases ${\bm{\theta}}^{H}$ through gradient descent, we alternatively choose to jointly optimize both the basis coefficients ${\bf w}$ and the adaptive bases ${\bm{\theta}}^{H}$ during the gradient descent step. During the least squares step, we solve a similar problem to equation 24 modified to suit the Fourier PINN architecture. Specifically, we solve the least squares problem of the form,
$\operatorname*{arg\,min}\nolimits_{\bf w}\|A^{\prime}{\bf w}-\textbf{b}\|_{\ell_{2}}^{2},$ (25)
such that $\textbf{b}_{i}={\mathcal{L}}[u_{F}]({\bf x}_{i})$ for the data points $\{{\bf x}_{i}\}_{i=1}^{M}$, and
@@
We define equation 25 for one operator ${\mathcal{L}}$, but the method extends to multiple operators. Additionally, we extend our least-squares method to non-linear operators, unlike many works that focus only on linear cases.
Specifically, when the operator ${\mathcal{L}}$ is nonlinear, the coefficient updates of ${\bf w}$ require more careful consideration. We consider non-linear operators ${\mathcal{L}}$, that can be decomposed into linear and nonlinear parts such as,
@@
where ${\mathcal{G}}$ is a linear operator and ${\mathcal{S}}$ is the nonlinear part. For illustration, consider the Allen-Cahn problem define by equation 14. We have ${\mathcal{G}}[u]=u_{xx}$ and ${\mathcal{S}}[u]=u(u^{2}-1)$, and solve equation 25 such that $\textbf{b}_{i}={\mathcal{L}}[u_{F}]({\bf x}_{i})-\textbf{s}_{i}$ for the data points $\{{\bf x}_{i}\}_{i=1}^{M}$. Here, $\textbf{s}_{i}={\mathcal{S}}[u_{F}]({\bf x})=u_{F}({\bf x}_{i};{\bf w})(u_{F}({\bf x}_{i};{\bf w})^{2}-1)$ are the Fourier PINN’s evaluations with the current ${\bf w}$ on the set of points, and
@@
In this case, since b is determined from the current ${\bf w}$ but fixed as a constant in the least squares estimation, the update of ${\bf w}$ is in essense a type of fixed point iteration. When ${\mathcal{L}}$ does not include any linear operator (which might be rare), we can use any continues optimization algorithm, e.g., L-BFGS, to update ${\bf w}$.
We integrate a custom basis pruning routine into the optimization algorithm to eliminate insignificant bases and refine model training. Our modified loss function defined in equation 23 enhances this hybrid optimization algorithm with an $L_{2}$ regularization term to identify and prioritize significant frequencies. Algorithm 1 outlines our adaptive learning routine. The algorithm begins with a joint optimization of ${\bm{\theta}}^{H}$ and ${\bf w}$ to provide a warm start to the model parameters. Next, we fix ${\bm{\theta}}^{H}$ while alternating between solving for the coefficients ${\bf w}$ using the least squares method and truncating bases with coefficients below a predefined threshold. Specifically, bases corresponding to $|w_{j}|\leq 10^{-3}$ are removed along with $w_{j}$. After completing the alternating solve and truncate steps for a set number of iterations, the algorithm jointly optimizes all remaining parameters and bases using the Adam optimization algorithm. We repeat this routine for a fixed number of iterations, followed by L-BFGS optimization until the final convergence.

### v2

@@ 4 Fourier Analysis of Strong BC PINNs
Both NNs and PINNs are known to exhibit spectral bias, meaning they readily capture low-frequency components of a target function but struggle with high-frequency details (Basri et al., 2020; Rahaman et al., 2019; Xu et al., 2020; Wang et al., 2021c). As demonstrated earlier, this bias causes the solution quality of PINNs to degrade as the target solution’s frequency increases. However, strong BC PINNs appear to mitigate or delay this issue, achieving better performance for higher-frequency solutions.
This section uses Fourier analysis to investigate why strong BC PINNs outperform standard PINNs in addressing spectral bias. The infinite Fourier series expresses a periodic function $u(x)$ with period $L$ as:
$u_{\infty}(x)=\sum_{n=-\infty}^{\infty}\widehat{u}[n]e^{i2\pi nx/L},$ (16)
where the Fourier coefficients are given by:
[…]
4.1 Numerical Fourier Analysis of Strong BC PINNs
To analyze spectral bias, we compute the Fourier coefficients for several manufactured solutions: 1. A single sine wave to demonstrate the baseline frequency handling of PINNs. 2. A combination of sine waves with differing frequencies to explore how multi-scale components are captured. 3. A Gaussian-modulated sine wave to evaluate performance on solutions with continuous spectral content. 4. A sine wave and polynomial combination to examine behavior under non-homogeneous boundary conditions.
For standard PINNs, the frequency spectrum typically shows a pronounced low-frequency peak, reflecting effective learning of low-frequency components and a gradual decay for higher frequencies, often resulting in heavy tails. These heavy tails indicate residual noise and the network’s difficulty in accurately filtering out high-frequency inaccuracies. By contrast, as we will show, strong BC PINNs exhibit faster decay in the higher-frequency components, enabling better suppression of noise and improved representation of fine details in the solution.
We trained a standard PINN and a strong BC PINN for each example with four hidden layers, 100 nodes per layer, and Tanh activations. Table 4 lists the specific training and hyperparameter details. We then computed the Discrete Fourier Transform (DFT) of the learned solutions $u_{\theta}(x)$, which represents their frequency spectrum. For $N$ equally spaced points $x_{j}$ sampled over the domain $[0,2\pi]$, the DFT is defined as:
$\widehat{u}_{\theta}[k]=\sum_{n=0}^{N-1}u_{\theta}(x_{n})e^{-i2\pi kn/N},\quad k=-\frac{N}{2},\dots,\frac{N}{2}-1.$
@@ 4.1 Numerical Fourier Analysis of Strong BC PINNs
Figure 3: Frequency spectrum of the learned solution (left) and the convolution operation (right) in the strong BC PINN and standard PINN. The ground-truth solution is $\sin(kx)$ with $k=15$. The left graph shows frequency handling by standard and strong BC PINNs, while the right demonstrates how convolution reduces high-frequency noise.
Figure 3 (left) compares the frequency spectrum of a standard PINN and a strong BC PINN trained on the 1D Poisson problem with the ground-truth solution $u(x)=\sin(kx)$, for $x\in[0,2\pi]$ and $k=15$. While the standard PINN successfully captures the target frequency, it also retains significant high-frequency noise, as evident in the heavy tails of the spectrum. In contrast, the strong BC PINN exhibits much faster decay in higher frequencies, effectively reducing high-frequency noise and improving accuracy. This indicates that the strong BC PINN filters out unnecessary components more efficiently, enhancing the clarity and precision of the learned solution by isolating the dominant frequency.
Figure 4: The frequency spectrum of the learned solutions (top) and absolute error of the Fourier coefficients (bottom) for the strong BC PINN and standard PINN compared to the ground truth. The ground-truth solution is $u(x)=\sin(2x)+\sin(16x)$.
Building on the analysis of single-frequency solutions in the previous section, we now examine the performance of standard and strong BC PINNs on a multi-scale target solution where multiple frequencies are present simultaneously. This setup provides a more challenging scenario, testing each model’s ability to balance between capturing dominant low-frequency components and accurately representing higher-frequency details. For a multi-scale target solution, $u(x)=\sin(2x)+\sin(16x)$, Figure 4 illustrates the frequency spectrum of the learned solutions. To simplify the analysis and focus on the unique information, we transition to visualizing only the positive frequency components, as the Fourier coefficients are symmetric for real-valued functions.
While the standard PINN captures the primary low-frequency components, it struggles to filter the higher-frequency components effectively, allowing residual high-frequency noise to persist. The strong BC PINN, however, captures both frequencies more accurately and suppresses extraneous high-frequency noise, leading to a lower overall error. This result highlights the strong BC PINN’s advantage in multi-scale problems, where it can effectively balance between capturing necessary frequencies and eliminating noise.
Figure 5: The frequency spectrum of the learned solutions (top) and absolute error of the Fourier coefficients (bottom) for the strong BC PINN and standard PINN compared to the ground truth. The ground-truth solution is $u(x)=e^{-0.5x^{2}}\sin(16x)$.
To evaluate the strong BC PINN’s performance on solutions with continuous spectral content, we used a Gaussian-modulated sinusoid, $u(x)=e^{-0.5x^{2}}\sin(16x)$. Figure 5 shows the Fourier spectrum for this case. The Gaussian modulation introduces a broad spectrum of frequencies, challenging the model’s ability to suppress irrelevant high frequencies while accurately capturing the primary frequency content. In this example, the strong BC PINN demonstrates robust noise reduction by filtering unnecessary frequencies, including low frequencies. This capability is essential for handling continuous spectra. The strong BC PINN can generalize well to functions with non-discrete frequency components, making it applicable to a broader range of real-world problems.
Figure 6: The frequency spectrum of the learned solutions (top) and absolute error of the Fourier coefficients (bottom) for the strong BC PINN and standard PINN compared to the ground truth. The ground-truth solution is $u(x)=x^{2}+\sin(16x)$.
Finally, we explore the strong BC PINN’s performance on non-homogeneous boundary conditions using the manufactured solution $u(x)=x^{2}+\sin(16x)$. Figure 6 illustrates the Fourier spectrum for this non-homogeneous boundary case. The polynomial term $x^{2}$ introduces a low-frequency component that is not harmonically related to the oscillatory term $\sin(16x)$. The standard PINN struggles to capture this low-frequency component fully. At the same time, the strong BC PINN demonstrates improved performance in accurately representing both the polynomial boundary effect and the oscillatory component. This example underscores the strong BC PINN’s effectiveness in handling non-homogeneous boundaries, as it can accommodate mixed low- and high-frequency components in the solution.
These results collectively illustrate the strong BC PINN’s enhanced ability to manage various frequency cases, from simple single-frequency cases to continuous spectra and complex non-homogeneous boundary conditions. The strong BC PINN demonstrates robustness and flexibility by effectively suppressing high-frequency noise and maintaining accuracy across various spectral components, making it suitable for a wide array of boundary-conditioned problems in applied settings.
4.2 Analytical Fourier Analysis of Strong BC PINNs
To analyze why multiplying the distance function ($\phi(x)$) with the NN in Equation 8 helps to obtain more accurate coefficients for high frequencies, we first represent the general distance function $\phi(x)$ as an infinite Fourier series in its discrete form [^1] While the analysis here is presented in the discrete Fourier series form for clarity and consistency with the numerical implementation, the results can be extended to the continuous Fourier transform by replacing summations over discrete indices $n$ with integrals over the continuous frequency domain. as in Equation 16,
@@ 4.2 Analytical Fourier Analysis of Strong BC PINNs
with period $L=2\pi$. We first analyze the polynomial distance function by substituting Equation 9 for $\hat{\phi}$ in Equation 18. From Equation 17, we get the Fourier coefficients,
@@
Similarly, to compare against our proposed adaptive distance function, we substitute in Equation 10 for $\hat{\phi}$ in Equation 18 and obtain the Fourier coefficients,
@@
Here, we expand the boundary functions $\phi(x)$ to be periodic (i.e., duplicating its definition in $[0,2\pi]$ to other intervals), and the period is $2\pi$. We can now obtain the Fourier transforms of each $\phi_{\infty}(x)$ by evaluating,
@@
Since $|\hat{\phi}_{poly}[n]|\propto 1/n^{2}$, the amplitude of the frequencies of $\phi_{poly}(x)$ decay quadratically fast with the increase of the absolute frequency value. This rapid decay is beneficial for denoising as it effectively suppresses high-frequencies, leading to smoother approximations. However, it may result in the loss of fine details, especially if essential signal components are present at higher frequencies. Alternatively, $|\hat{\phi}_{exp}[n]|\propto\alpha/(\alpha^{2}+n^{2})$, indicating that the frequency amplitudes follow Lorentzian decay controlled by the decay rate $\alpha$. Lorentzian decay is less aggressive in reducing high-frequency components, which helps preserve fine details and sharp transitions in the signal. This flexibility helps tune the decay based on signal characteristics. However, the slower decay rate may lead to less effective noise suppression.
@@
$\hat{u}_{\theta}(\omega)=\hat{\phi}*\widehat{u}_{N}=\int_{-\infty}^{\infty}\hat{\phi}(\omega-\nu)\;\widehat{u}_{N}(\nu)\;{\rm d}\nu$ (22)
where * denotes convolution, and $\widehat{u}_{N}$ is the Fourier transform of the NN. Suppose $\omega>0$, we know that $\hat{\phi}(\omega-\nu)$ is obtained by first reflecting $\hat{\phi}(\nu)$ about the y-axis, and then shifting the frequency $\omega$ left along the spectrum. Then, we integrate $\widehat{u}_{N}(\nu)$ weighted by $\hat{\phi}(\omega-\nu)$. The larger the frequency $\omega$, the more $\hat{\phi}$ is moved left to obtain $\hat{\phi}(\omega-\nu)$ resulting in a stronger weighting of the high-frequency components during integration. Since the frequency coefficients of $\hat{\phi}$ decay quadratically fast ($|\hat{\phi}[n]|\propto 1/n^{2}$), the larger the portion of the tail is used, and the smaller the integration result. Specifically, with $\phi$, the corresponding amplitude of $\hat{u}_{\theta}(\omega)$ decrease fast when $\omega$ increases [^2] The same conclusion applies when we consider $\omega<0$.. See Figure 3 for an illustration in the discrete case. During training, the boundary function pushes the surrogate model in Equation 8 to weaken irrelevant high-frequency components to reduce large tails and better capture the frequency spectrum.
@@ 5 Fourier PINNs
Our Fourier analysis in the previous section revealed that the standard PINNs inefficiently capture the true amplitude of frequencies in the power spectrum of their predicted solutions. Specifically, the analysis showed that although standard PINNs could identify the true solution frequency, they failed to prune large and noisy frequencies, as indicated by the large amplitudes assigned at the tail end of the power spectrum in Figure 3. While the strong BC PINNs address this issue by multiplying the NN with a boundary function—which helps decay high-frequency amplitudes faster—their overall benefit remains limited. As shown in Figure 2, performance degrades as $k$ increases, indicating the boundary function’s inability to learn ultra-high frequency solutions. The strong BC PINNs are also restricted to specific boundary conditions, such as Dirichlet conditions in Equation 7, limiting their applicability.
@@ 5.1 Fourier PINNs
We introduce Fourier PINNs, a novel PINN architecture to overcome these challenges. Fourier PINNs seamlessly handle diverse boundary conditions, like the standard PINNs, while emphasizing the true solution frequencies during training—akin to the strong BC PINNs. Specifically, in order to flexibly yet comprehensively capture the frequency spectrum, we first introduce a set of dense frequency candidates, $\{\omega_{n}\}_{n=1}^{K}$, evenly sampled from the range $[1,K]$. From this set of frequency candidates, we define a set of trainable Fourier bases $u_{B}$ as,
@@
While we present our method in the one-dimensional input case, extending it to multiple dimensions is straightforward. Consider the two-dimensional input ${\bf x}=[x_{1},x_{2}]$ for an example. We use the same set of frequencies to construct Fourier bases for each input. Specifically, we construct
@@
Then, we apply the cross-product of $\bm{\phi}(x_{1})$ and $\bm{\phi}(x_{2})$ to obtain the tensor-product Fourier bases for ${\bf x}$. The surrogate model is given by,
@@ 5.2 Adaptive Basis Selection Training Algorithm
To aid in identifying the significant frequencies, we add an $L_{2}$ regularization term to the basis parameters ${\bf w}$. While $L_{2}$ regularization does not promote sparsity, in our specific case, we use $L_{2}$ regularization with a pruning strategy, which indirectly promotes sparsity by pruning coefficients after driving them toward a threshold. Methods such as the SINDy (Sparse Identification of Nonlinear Dynamics) algorithm Zhang & Schaeffer (2019) have explored this strategy, where $L_{2}$ regularization helps to stabilize the coefficients during optimization before pruning—yielding sparse solutions. This combination effectively balances stability and sparsity, as shown in previous work Wang et al. (2021a). In this case, we chose not to use $L_{1}$ regularization due to practical challenges, including the instability it can introduce during training, particularly when optimizing neural networks for complex systems. $L_{2}$ regularization, combined with a careful pruning process, offers more stable convergence while still leading to sparse solutions after pruning the small coefficients.
This additional regularization penalizes large coefficients for the basis functions, thereby promoting sparsity and encouraging the model to focus on learning the most significant frequencies. The full loss function we use is similar to PINNs (see Equation 4) but with the addition of the $L_{2}$ regularization term. Specifically, the loss function is,
@@
where $\|{\bf w}\|^{2}$ represents the $L_{2}$ regularization of ${\bf w}$ and $\alpha$ is the regularization strength defined by the user before initiating training. This regularization term helps prune or inhibit unnecessary frequencies, allowing the model to focus more on learning and improving the bases that significantly contribute to the solution.
Linear Operators
Our adaptive basis selection algorithm optimizes the regularized loss function in Equation 26 by building upon the hybrid least squares gradient descent method proposed by Cyr et al. (2020). Cyr et al. (2020) formulates the output of a neural network as a linear combination of nonlinear basis functions, as shown in Equation 3. Their algorithm alternates between optimizing the hidden weights (${\bm{\theta}}^{H}$) via gradient descent and the final layer coefficients ($\mathbf{c}$) of the NN through the least squares problem of the form,
$\operatorname*{arg\,min}\nolimits_{\textbf{c}}\|A\textbf{c}-\textbf{y}\|_{\ell_{2}}^{2},$ (27)
such that $\textbf{y}_{i}={\mathcal{L}}[u]({\bf x}_{i})$ for the data points $\{{\bf x}_{i}\}_{i=1}^{M}$, $A_{ij}={\mathcal{L}}[\psi_{j}({\bf x}_{i};\;{\bm{\theta}}^{H})]$ for $j=1,...,W$, and ${\mathcal{L}}$ is some linear operator. Equation 27 extends to problems with multiple operators, such as those defined by Equation 1 and Equation 2 when ${\mathcal{F}}$ and ${\mathcal{B}}$ are linear operators.
Representing the output of Fourier PINNs in Equation 23 as a linear combination of both adaptive bases (i.e., the hidden layers of the NN) and an augmented Fourier series ensures our model’s compatibility with the general approach of the alternating least squares and gradient descent optimization. While Cyr et al. (2020) hold the coefficients of the last layer in the NN (c) constant while optimizing the bases ${\bm{\theta}}^{H}$ through gradient descent, we alternatively choose to jointly optimize both the Fourier PINN’s basis coefficients ${\bf w}$ (which are comprised of both the NN basis coefficients c along with the Fourier basis coefficients $a_{1},...,a_{K}$ and $b_{1},...,b_{K}$) and the adaptive bases ${\bm{\theta}}^{H}$ during the gradient descent step. During the least squares step, we solve a similar problem to Equation 27 modified to suit the Fourier PINN architecture. Specifically, we solve the least squares problem of the form,
$\operatorname*{arg\,min}\nolimits_{\bf w}\|A^{\prime}{\bf w}-\textbf{y}\|_{\ell_{2}}^{2},$ (28)
such that $\textbf{y}_{i}={\mathcal{L}}[u]({\bf x}_{i})$ for the data points $\{{\bf x}_{i}\}_{i=1}^{M}$, and
@@
We define Equation 28 for one operator ${\mathcal{L}}$, but the method extends to multiple operators. Additionally, we extend our least-squares method to nonlinear operators, unlike many works that focus only on linear cases.
Non-Linear Operators
When the operator ${\mathcal{L}}$ is nonlinear, updating the coefficients ${\bf w}$ requires additional care to ensure the method accurately handles the nonlinearities and achieves convergence. Nonlinear operators can often be decomposed into a combination of linear and nonlinear components, simplifying the process of updating ${\bf w}$ during optimization. Specifically, we decompose the operator ${\mathcal{L}}$ as follows:
@@
where ${\mathcal{G}}$ represents the linear part of the operator, and ${\mathcal{S}}$ denotes the nonlinear component.
To illustrate this, consider the Allen-Cahn equation, where the operator ${\mathcal{L}}$ contains both linear and nonlinear terms. In this case, the linear part ${\mathcal{G}}$ is $u_{xx}$, and the nonlinear term ${\mathcal{S}}[u]$ is $u(u^{2}-1)$, leading to the decomposition:
${\mathcal{G}}[u]=u_{xx},\quad{\mathcal{S}}[u]=u(u^{2}-1).$ (31)
We solve this system by incorporating the Fourier PINN formulation. For the least squares solution of the system, the nonlinear operator ${\mathcal{S}}$ is applied at each iteration using the current approximation of $u$, which is based on the current coefficients ${\bf w}$ and is given as,
$\textbf{s}_{i}={\mathcal{S}}[u_{F}]({\bf x}_{i})=u_{F}({\bf x}_{i};{\bf w})(u_{F}({\bf x}_{i};{\bf w})^{2}-1)$ (32)
Here, $u_{F}({\bf x}_{i};{\bf w})$ is the Fourier PINN’s evaluation at point ${\bf x}_{i}$, using the current parameters ${\bf w}$. The linear part ${\mathcal{G}}$ is handled directly in the least squares system, and the matrix $A^{\prime}$ is defined as:
@@
The nonlinear terms $\textbf{s}_{i}$ are included in the target vector such that $\textbf{y}_{i}={\mathcal{L}}[u]({\bf x}_{i})-\textbf{s}_{i}$, which is updated iteratively and reduces the problem to a linear least-squares system. As this method incorporates the current values of ${\bf w}$ into calculating the nonlinear components, the optimization process can be viewed as a form of fixed-point iteration. We note that for cases where ${\mathcal{L}}$ consists primarily of nonlinear terms or does not include a linear operator, the optimization problem might need to be handled using continuous optimization algorithms such as L-BFGS, which effectively updates ${\bf w}$ based on the entire system to ensure convergence.
We integrate a custom basis pruning routine into the optimization algorithm to eliminate insignificant bases and refine model training. Our modified loss function defined in Equation 26 enhances this hybrid optimization algorithm with an $L_{2}$ regularization term to identify and prioritize significant frequencies. Algorithm 1 outlines our adaptive learning routine. The algorithm begins with a joint optimization of ${\bm{\theta}}^{H}$ and ${\bf w}$ to provide a warm start to the model parameters. Next, we fix ${\bm{\theta}}^{H}$ while alternating between solving for the coefficients ${\bf w}$ using the least squares method and truncating bases with coefficients below a predefined threshold. Specifically, bases corresponding to $|w_{j}|\leq 10^{-3}$ are removed along with $w_{j}$. After completing the alternating solve and truncating steps for a set number of iterations, the algorithm jointly optimizes all remaining parameters and bases using the Adam optimization algorithm. We repeat this routine for a fixed number of iterations, followed by L-BFGS optimization until the final convergence.